### v1

# Marginal and conditional summary measures: transportability and compatibility

Abstract
Marginal and conditional summary measures do not generally coincide, have different interpretations and correspond to different decision questions. While these aspects have primarily been recognized for non-collapsible summary measures, they are also problematic for some collapsible measures in the presence of effect modification. We clarify the interpretation and properties of different marginal and conditional summary measures, considering different types of outcomes and hypothetical outcome-generating mechanisms. We describe implications of the choice of summary measure for transportability, highlighting that covariates not conventionally described as effect modifiers can modify population-level treatment effects. Finally, we illustrate existing summary measure incompatibility issues in the context of evidence synthesis, using the case of covariate adjustment methods for indirect treatment comparisons. Because marginal and conditional summary measures do not generally coincide, their naïve pooling in evidence synthesis can produce bias. Almost invariably, care is needed to ensure that evidence synthesis methods are combining compatible summary measures, and this may be easier to ensure with full access to individual patient data.
keywords:
estimands; evidence synthesis; summary measure; meta-analysis; transportability; indirect treatment comparisons
1 Introduction
Randomized controlled trials (RCTs) are the “gold standard” research design for estimating causal treatment effects with high internal validity. Randomization ensures that treatment assignment is independent of all measured and unmeasured baseline covariates, thereby guaranteeing that these are balanced in expectation over the trial’s treatment arms. Nevertheless, high internal validity does not necessarily guarantee external validity with respect to a target population of substantive interest. External validity is threatened by differences in the distribution of effect modifiers – covariates inducing variation in treatment effects – between the RCT and the target population (Westreich et al., 2019).
Differences in the distribution of effect modifiers can compromise the transportability or constancy of treatment effects across populations (stat03728, 2014; stat05152, 2015). While said differences are recognized as a fundamental component of treatment effect heterogeneity, a lesser-known source of heterogeneity is the summary (effect) measure used to compare outcomes between treatment arms. Summary measures can be marginal (unconditional) or conditional; directly collapsible (mean and risk difference), collapsible but not directly collapsible (risk ratio) or non-collapsible (odds ratio and hazard ratio) (stat05130, 2015). Definitions of these terms are provided in Table 1.
Marginal and conditional summary measures do not generally coincide and correspond to different decision questions, with discussion over which is more suitable for healthcare decision-making (Phillippo et al., 2021; Remiro-Azócar, 2022b; Phillippo et al., 2025). Whether a summary measure is marginal or conditional has implications on its transportability across populations. In evidence synthesis, the naïve pooling of marginal and conditional measures can produce bias (Remiro-Azócar et al., 2021b, 2022), particularly for analyses using aggregate-level data, as trials may report different summary measures in publications. This issue of “estimand incompatibility” has primarily been recognized for non-collapsible summary measures, where marginal and conditional measures do not coincide, even in the absence of effect modification by the covariates (Daniel et al., 2021). Nevertheless, the issue is also problematic for some collapsible measures – those that are not directly collapsible – in the presence of effect modification (Remiro-Azócar, 2024).
In this article, we will clarify the interpretation of different marginal and conditional summary measures. We do not intend to express a preference towards one or another, but to set out the relevant terminology and illustrate the properties of relevant summary measures. We consider different types of outcomes and outcome-generating mechanisms, with and without effect modification. We describe implications of the choice of summary measure for transportability, highlighting that covariates not conventionally described as effect modifiers can modify population-level treatment effects. We conclude by illustrating existing summary measure incompatibility issues in the context of evidence synthesis, using the case of covariate adjustment methods for indirect treatment comparisons. Almost invariably, care is needed to ensure that evidence synthesis methods are combining compatible summary measures.
Table 1: Definitions of key terminology for summary measures.
Term Definition
Collapsible
Summary measure for which the marginal measure can always be expressed as a weighted average of individual-
or subgroup-level conditional measures
Conditional
Summary measure contrasting (functions of) the conditional outcome distributions under different treatments,
given specific covariate values
Directly collapsible
Collapsible summary measure for which the collapsibility weights are given by the marginal distributions of the
covariates that are conditioned on, e.g., proportions for discrete covariates (Colnet et al., 2023)
Marginal
Summary measure contrasting (functions of) the marginal or unconditional outcome distributions under
different treatments
Non-collapsible
Summary measure for which the marginal measure does not equal a weighted average of conditional measures,
even where the conditional measure is equal across all individuals or subgroups (Huitfeldt et al., 2019)
2 Marginal and conditional estimands
An estimand, as defined by the International Council of Harmonisation (ICH) E9 (R1) Addendum, is a precise definition of the treatment effect that is targeted by a clinical trial (Keene et al., 2023). The notion of an estimand as a target of estimation is not new, but its application to pose research questions in registrational RCTs has been encouraged by ICH E9 (R1). The summary measure is only one of the five attributes of estimands, as per ICH E9 (R1). Nevertheless, because our focus is on said attribute, we will use the terms summary measure and estimand interchangeably. Following the potential outcomes framework, these are defined as causal contrasts of any function of the distributions of counterfactual outcomes under different treatments for the same set of subjects (Hernán & Robins, 2020). Because ICH E9 (R1) is aligned with counterfactual language, potential outcomes notation is useful to formally construct the summary measures.
Consider a hypothetical ideal RCT targeting a population of substantive interest. Let $Y^{t}$ denote the potential outcome that might have been observed for an individual assigned to binary treatment $T=t$, with $t\in\{0,1\}$ indicating assignment to the control and the active treatment group, respectively. Each participant in the trial has two potential outcomes, $Y^{1}$ and $Y^{0}$. Assuming no loss to follow up and complete data, one of these is observed in the trial. The other is the outcome hypothetically realized under a different treatment than that actually assigned.
The marginal treatment effect (MTE) estimand is a contrast of functions of marginal or unconditional outcome expectations under each treatment. On the additive scale:
$\textnormal{MTE}=g\left(E\left(Y^{1}\right)\right)-g\left(E\left(Y^{0}\right)\right),$ (1)
where $g(\cdot)$ is a suitable “link” function, transforming the potential outcome means onto the minus/plus infinity range, and each expectation $E(\cdot)$ is over the distribution of potential outcomes in the trial population. The marginal estimand is typically described as the average treatment effect in the trial population, had everyone in the trial population been assigned active treatment versus control (Austin, 2014). Note that the use of a link function is not strictly necessary for the definition and is usually obviated in the causal inference literature.
The marginal estimand is often considered of primary interest for population-level policy decisions (Remiro-Azócar, 2022b). As per Equation 1, the causal inference literature tends to define population effects through functions of the marginal distributions of outcomes, although “the word marginal is a preference rather than necessary to the definition” because marginal estimands appear to be favored for causal inference based on observational data (Morris et al., 2022). Consequently, the marginal treatment effect is commonly referred to as the average treatment effect (ATE) or the population-average treatment effect (PATE) (Austin, 2013; Remiro-Azócar, 2024), but we will not use this terminology to avoid confusion with other (population-)average treatment effects introduced later in this section.
Conditional treatment effect estimands contrast functions of outcome expectations that are conditional on baseline covariates, with these playing an explicit role in the treatment effect definition. There are different conditional estimands that are potentially of interest to researchers, with different interpretations and corresponding to different decision problems. We make a conceptual distinction between three conditional summary measures: the conditional treatment effect evaluated at particular covariate values, the conditional treatment effect at the (covariate) means, and the population-average conditional treatment effect.
While, by virtue of randomization, the marginal estimand within the ideal RCT can be identified from the observed data with minimal assumptions, the identification of conditional estimands within such RCT may require statistical assumptions about correct model specification, particularly with continuous covariates or where there are only a few subjects in given covariate strata (Van Lancker et al., 2022, 2024).
We define the conditional treatment effect (CTEX) given a particular covariate value $X=x$ as:
$\textnormal{CTEX}=g\left(E\left(Y^{1}\mid X=x\right)\right)-g\left(E\left(Y^{0}\mid X=x\right)\right),$ (2)
where each expectation $E(\cdot)$ is over the distribution of potential outcomes in the trial population, but conditional on covariate values. The CTEX is the average treatment effect had a subset of individuals, with the same covariate profile, been assigned active treatment versus control. It is often referred to as the conditional average treatment effect (CATE) (Colnet et al., 2023), but we will not use this term to avoid confusion with other average conditional effects introduced later in this section.
To simplify the exposition, we have assumed that there is a single baseline covariate. If this is binary or categorical, the CTEX can be interpreted as a subgroup- or stratum-specific effect. If multiple relevant covariates are conditioned on, it comes closer to an individualized or subject-specific effect. The CTEX is of interest to clinicians and individual patients, particularly if subjects with different covariate values benefit differently from treatment. It is arguably of lesser interest for population-level decision-making, especially if the treatment effect does not change with the covariate values; but can be useful when considering decisions for different subgroups or subpopulations if the treatment effect does change.
The conditional treatment effect at the (covariate) means (CTEM) is a special case of the CTEX:
$\textnormal{CTEM}=g\left(E\left(Y^{1}\mid X=\bar{X}\right)\right)-g\left(E\left(Y^{0}\mid X=\bar{X}\right)\right),$ (3)
with $\bar{X}=E(X)$. This reflects the conditional treatment effect for subjects with covariate values equal to their mean. This summary measure may be of interest if the covariate is continuous. Nevertheless, it is not generally a desirable target for population-level decision-making and likely not meaningful for discrete covariates.
Consider sex, ethnicity, smoking status, biomarker status or exposure to previous treatment, where the mean is the proportion of subjects in a category. The CTEM does not make much sense and may have an awkward or confusing interpretation, denoting the effect for a non-existent “average subject” in the middle of the categories. Confusingly, the CTEM and conditional outcome expectations $E\left(Y^{t}\mid X=\bar{X}\right)$ for $t\in\{0,1\}$ are sometimes known as marginal effects (at the mean) and marginal (or “least-squares”) means, respectively, in the econometrics literature (Arel-Bundock et al., 2024; Lenth, 2016).
Finally, the population-average conditional treatment effect (PACTE) is defined as:
PACTE $=E_{X}\left(\textnormal{CTEX}\right)$
$=E_{X}\left(g\left(E\left(Y^{1}\mid X=x\right)\right)-g\left(E\left(Y^{0}\mid X=x\right)\right)\right),$ (4)
where each inner expectation $E(\cdot)$ is over the distribution of potential outcomes in the trial population, conditional on covariate values, and the outer expectation $E_{X}(\cdot)$ is over the trial population. Namely, the PACTE is the average CTEX across all the subjects or subgroups in the trial population. While the terms population-average and marginal are often used interchangeably, the PACTE is also a population-average summary measure because it applies to the whole trial population and is an average over the covariate distribution of such population.
It has been argued that the PACTE can be interpreted as the conditional treatment effect at the (covariate) means (Remiro-Azócar et al., 2021a). Nevertheless, as we shall discuss in Section 3, the PACTE only coincides with the CTEM where the CTEX is constant or a linear function of $X$. The PACTE and the CTEM are fundamentally distinct summary measures. Note that while the CTEX is a function of the covariate(s), the CTEM and the PACTE (and the MTE) compress the treatment effect surface into a single aggregate summary.
The difference between the MTE and the PACTE is the order of operations, or the scale on which the averaging takes place (Phillippo et al., 2025). The MTE takes the unconditional expectation of the potential outcomes on their natural scale, then contrasts the transformed averages on the scale imposed by the link function:
MTE $=g\left(E\left(Y^{1}\right)\right)-g\left(E\left(Y^{0}\right)\right)$
$=E_{X}\left(g\left(E\left(Y^{1}\mid X=x\right)\right)\right)-E_{X}\left(g\left(E\left(Y^{0}\mid X=x\right)\right)\right),$
by the law of total expectation. Conversely, the PACTE contrasts the conditional expectations of the potential outcomes on the scale imposed by the link function, then takes the average of the contrasts, with the outer expectation over $X$ outside the link functions. The changing order of operations is irrelevant where $g(\cdot)$ is the identity link, but otherwise matters.
3 Model-based estimands
The summary measure definitions in Equations 1 to 4 have been formulated on an additive scale. Such definitions are “model-free”; despite the use of function $g(\cdot)$, the estimands do not necessarily rely on modeling assumptions and are not necessarily encoded by the coefficients of a model (Van Lancker et al., 2024; Vansteelandt & Dukes, 2022; Mütze et al., 2025). Nevertheless, in practice, the choices of $g(\cdot)$ and the scale of the summary measure are often “model-based”: influenced by modeling preferences or statistical considerations for different outcome types (Van Lancker et al., 2024; Daniel et al., 2021; Mütze et al., 2025).
For instance, the logit link function is used to model binary outcomes because fitted logistic regressions predict outcome probabilities between zero and one. For non-negative integer (“count”) outcomes, the logarithmic link naturally incorporates the lower bound of a count at zero. Typically, within a generalized linear modeling (GLM) framework: for continuous, count or binary outcomes, the link function $g(\cdot)$ is a canonical link; the identity, logarithmic or logit link, respectively, with the mean difference, log rate ratio or log odds ratio scale, respectively, as the additive linear predictor scale imposed by the link function of the GLM (Wei et al., 2024; Højbjerre-Frandsen et al., 2025).
Admittedly, investigators are free to derive treatment effects on other scales – e.g., one can use a logit link function to model binary outcomes with a logistic regression, then report the treatment effect as a difference of (marginal) risks, using an identity link for summarization. Nevertheless, as is often the case in practice for RCTs (Wei et al., 2024; Mütze et al., 2025), we shall assume that the scale for the target summary measure is determined by the GLM canonical link function that would be used to model the outcome type. Additionally, we suppose that the conditional outcome expectations follow concrete parameterizations with specific functional forms about the effect of covariates on conditional measures. This implies that the conditional estimands are known and can be expressed in simple terms composed of model parameter(s). In reality, the “true” generative function for the conditional outcome expectation will be more complex, and one is likely to have incomplete knowledge on functional forms for the baseline covariates.
Under the GLM framework, we postulate three hypothetical parametric outcome-generating models to explore the behavior of estimands under different generative mechanisms. These are not exhaustive and have been posited for illustrative purposes. The first outcome-generating mechanism induces treatment effect homogeneity across subjects, such that the CTEX on the linear predictor scale is constant. In this scenario, the MTE, CTEM and PACTE are equivalent for collapsible summary measures, but the MTE does not coincide with any of the conditional quantities for non-collapsible summary measures.
The second outcome-generating mechanism induces treatment effect heterogeneity with the CTEX varying as a linear function of $X$. In this scenario, the MTE, CTEM and PACTE are equivalent for directly collapsible summary measures, but the MTE does not coincide with any of the conditional quantities for summary measures that are not directly collapsible. The third outcome-generating mechanism also induces treatment effect heterogeneity but with the CTEX varying as a non-linear – in this case, quadratic – function of $X$. This highlights a situation where the CTEM and the PACTE are not equivalent, regardless of whether there is collapsibility (or direct collapsibility) or not.
3.1 Homogeneous CTEX
The outcome model with treatment effect homogeneity, denoted the homogeneous illustrative model, specifies the conditional expectation of potential outcome $Y^{t}$ for treatment $T=t$ given $X$ as:
$E(Y^{t}\mid X)=g^{-1}\left(\beta_{0}+\beta_{X}X+\beta_{T}t\right),$ (5)
where $g(\cdot)$ is a suitable link function, $\beta_{0}\in\mathbb{R}$ is an intercept quantifying the “baseline risk” on the linear predictor scale (the conditional outcome expectation on such scale for $X=0$ and $T=0$), and the model coefficients $\beta_{X},\beta_{T}\in\mathbb{R}$ quantify conditional predictor-outcome associations. We place no further statistical assumptions, e.g., normality, on the conditional outcome distribution. Because the outcome model does not include a covariate-treatment product (“interaction”) term but does include a term for the main covariate effect, covariate $X$ is said to be purely prognostic (Ballman, 2015). The baseline risk $\beta_{0}$ is typically a proxy for unobserved prognostic factors.
For a given link function, substituting Equation 5 into Equation 2 yields $\textnormal{CTEX}=\beta_{T}$ on the linear predictor scale. In this case, the CTEX is not actually covariate-specific; it is identical across all individuals or subgroups regardless of their covariate value. As the CTEX does not vary with the level of $X$, the model in Equation 5 imposes treatment effect homogeneity on the linear predictor scale. The conditional treatment effect evaluated at the (covariate) mean $\textnormal{CTEM}=\beta_{T}$ is the same as for every other value of the covariate. Because the PACTE averages the constant CTEX across the subjects in the population, $\textnormal{PACTE}=\beta_{T}$.
With respect to the marginal estimand, consider that $g(\cdot)$ is the identity link, such that the outcome-generating model is linear. On the mean difference scale, by the law of total expectation, the marginal treatment effect estimand $\textnormal{MTE}=\beta_{T}$, and the model-based coefficient $\beta_{T}$ coincides with the model-free definition of the marginal estimand and all conditional estimands (Remiro-Azócar, 2024). With the homogeneous linear generative model, the marginal mean difference does not depend on the distribution of the purely prognostic covariate $X$.
Now, consider that $g(\cdot)=\log(\cdot)$, such that the outcome-generating mechanism is log-linear; for instance, a Poisson or negative binomial model. These models also add the logarithm of person-time, $\log(\tau)$, as an offset to the right-hand side of Equation 5. We shall assume that person-time is fixed across subjects, e.g., corresponding to the trial follow-up period, but we return to this caveat later in this section. On the log rate ratio scale, $\textnormal{MTE}=\beta_{T}$, and the model-based coefficient $\beta_{T}$ coincides with the model-free definition of the marginal estimand and all conditional estimands (Remiro-Azócar, 2024; Vellaisamy & Vijay, 2008). With the homogeneous log-linear generative model (and constant person-time), the marginal (log) rate ratio does not depend on the distribution of the purely prognostic covariate $X$.
Finally, consider that $g(\cdot)=\operatorname{logit}(\cdot)$, such that the outcome-generating model is logistic. Because the (log) odds ratio is non-collapsible (Daniel et al., 2021), the marginal summary measure cannot be expressed as a weighted average of individual- or subgroup-level CTEX, even in the absence of treatment effect heterogeneity (Colnet et al., 2023; Huitfeldt et al., 2019). Assuming $\beta_{X}\neq 0$, the marginal and conditional (log) odds ratios are, almost invariably, not equal, with the marginal estimand closer to the null hypothesis – zero on the log odds ratio scale and one on the odds ratio scale – than all the conditional estimands.
Via Jensen’s inequality and the monotonicity of the logit function, it can be shown that: (1) $\beta_{T}>\textnormal{MTE}$ if $\beta_{T}>0$; and (2) $\beta_{T}<\textnormal{MTE}$ if $\beta_{T}<0$ (Colnet et al., 2023). While the model-based coefficient $\beta_{T}$ can be interpreted as a conditional estimand, it cannot be interpreted as a population-average marginal estimand because it is not equal to the model-free definition of the MTE, despite enforcing treatment effect homogeneity across subjects on the (log) odds ratio scale. Moreover, the MTE on such scale does not only depend on $\beta_{T}$, but depends on the full distribution of the purely prognostic covariate $X$, not just its mean (Daniel et al., 2021), as well as the distribution of baseline risk because prognostic factors may not even be known or measured (Phillippo et al., 2025).
Mean differences, risk differences and risk ratios are collapsible. One would think that (log) rate ratios inherit collapsibility from (log) risk ratios, as rates are risks per unit time. However, (log) rate ratios are non-collapsible (Sjölander et al., 2016), unless person-time is not influenced by the outcomes (treatment or baseline covariates), for instance, in a study of non-fatal recurrent events where all subjects remain at risk throughout the trial follow-up period. We will assume that this is the case for ease of exposition, something which allows us to describe (log) rate ratios as (log) risk ratios when discussing count outcomes and log-linear outcome models.
Table 2 presents a taxonomy of marginal estimands based on the homogeneous illustrative models. Figure 1 indicates whether different estimands match for the homogeneous illustrative models, across link functions and summary measures. The CTEX is not displayed because, outside the homogeneous illustrative model, it is a function of the covariates and not a “one number” summary. The shaded squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent by definition. In Table 2 and Figure 1, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
Table 2: Model-based marginal estimands for the homogeneous illustrative models. For count outcomes and the log link, person-time is assumed constant, such that the log rate ratio is collapsible and can be interpreted a log risk ratio.
Outcome Link function Summary measure Marginal estimand
Continuous Identity Mean difference Does not depend on the distribution of purely prognostic covariates
Count Logarithmic Log risk ratio Does not depend on the distribution of purely prognostic covariates
Binary Logit Log odds ratio Depends on the full joint distribution of purely prognostic covariates
{logicpuzzle}
[…]
(a) Identity link, mean difference
{logicpuzzle}
[…]
(b) Logarithmic link, log risk ratio
{logicpuzzle}
[…]
(c) Logistic link, log odds ratio
Figure 1: Matrices indicating whether different estimands are equivalent for the homogeneous illustrative models. The shaded squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
Outside the GLM framework, a model and summary measure that are widely used for the analysis of time-to-event outcomes in RCTs are the Cox proportional hazards model and the (log) hazard ratio. While the Cox model uses a logarithmic link to map hazards to the linear predictor, (log) hazard ratios are non-collapsible for similar reasons to (log) rate ratios, due to conditioning on past survival (Sjölander et al., 2016).
The marginal (log) hazard ratio depends on the shape of the baseline hazard and on the full joint distribution of purely prognostic covariates that do not “interact” with treatment, even under a generative model enforcing the constancy of the conditional (log) hazard ratio across covariate values (Martinussen & Vansteelandt, 2013; Daniel et al., 2021; Phillippo et al., 2025). Awkwardly, such generative model would result in non-proportional hazards for the marginal summary measure because proportional hazards cannot hold simultaneously on the marginal and conditional scales (Daniel et al., 2021; Phillippo et al., 2025).
In summary, while the within-trial estimation of the MTE generally requires weaker statistical assumptions than the estimation of the conditional summary measures, it likely requires stronger assumptions for transportability across different populations. This is particularly the case for non-collapsible summary measures, given their MTE dependence on the joint distribution of purely prognostic covariates, even in the absence of treatment effect heterogeneity.
3.2 Linear heterogeneous CTEX
The homogeneous illustrative model in Section 3.1 is sometimes considered too restrictive, in which case one can postulate an outcome model that accommodates for treatment effect heterogeneity across individuals. The heterogeneous illustrative model specifies the conditional expectation of potential outcome $Y^{t}$ for treatment $T=t$ given $X$ as:
$E(Y^{t}\mid X)=g^{-1}\left(\beta_{0}+\beta_{X}X+\beta_{T}t+\beta_{XT}Xt\right).$ (6)
Here, $X$ is prognostic of outcome, with the coefficient $\beta_{X}\in\mathbb{R}$ quantifying the conditional covariate-outcome association in the control group.
Assuming that the interaction coefficient $\beta_{XT}\in\mathbb{R}$ is non-null, $X$ also modifies the (conditional) treatment effect on the linear predictor scale. In this case, covariate $X$ is said to be an effect modifier. The RCT literature has typically conceptualized effect modifiers in this way, as baseline covariates that induce treatment effect heterogeneity through treatment-covariate interactions in an outcome-generating model (Christensen et al., 2021).
For a given link function, substituting Equation 6 into Equation 2 gives $\textnormal{CTEX}=\beta_{T}+\beta_{XT}x$ at $X=x$ on the linear predictor scale. That is, for the identity, logarithmic or logit link, the corresponding conditional mean difference, log risk ratio or log odds ratio, respectively, depends on the level of $X$. As such, $\beta_{T}$ is only the conditional treatment effect evaluated at $X=0$, and there is no longer a single CTEX estimand. For a given link function, substituting Equation 6 into Equation 3 yields $\textnormal{CTEM}=\beta_{T}+\beta_{XT}\bar{X}$. Similarly, substituting Equation 6 into Equation 4 gives $\textnormal{PACTE}=E_{X}\left(\beta_{T}+\beta_{XT}X\right)=\beta_{T}+\beta_{XT}E_{X}\left(X\right)=\beta_{T}+\beta_{XT}\bar{X}$.
With respect to the marginal estimand, consider that $g(\cdot)$ is the identity link, such that the outcome-generating model is linear. On the mean difference scale, by the law of total expectation, $\textnormal{MTE}=\beta_{T}+\beta_{XT}\bar{X}$ (Remiro-Azócar, 2024). While $\beta_{T}$ no longer has a marginal interpretation, the marginal estimand coincides with the CTEM and PACTE, and can still be expressed in terms of model coefficients and the unconditional expectation of $X$. This is because the mean difference is directly collapsible (Huitfeldt et al., 2019). Namely, the marginal measure is equal to a weighted average of conditional measures, with the weights given by marginal distributions of the covariates that are conditioned on, e.g., for binary or categorical covariates, the weights would be the covariate proportions (Colnet et al., 2023).
Unlike the mean difference, the (log) risk ratio is not directly collapsible (Kiefer & Mayer, 2019). The collapsibility weights for this summary measure are more complex; the marginal (log) risk ratio is not generally equal to a weighted average of conditional (log) risk ratios, if the weights are given by marginal covariate summary moments (Huitfeldt et al., 2019). Consequently, with $g(\cdot)=\log(\cdot)$ and the heterogeneous log-linear generative model, the expression for the marginal (log) risk ratio cannot simply be reduced to one involving $\beta_{T}$, $\beta_{XT}$ and $\bar{X}$. In this scenario, the MTE on the log risk ratio scale is not equivalent to the CTEM or the PACTE (Remiro-Azócar, 2024).
Moreover, where there are multiple prognostic baseline covariates and in the presence of treatment effect heterogeneity, the marginal (log) risk ratio cannot be identified using the marginal distributions of the effect modifiers (Remiro-Azócar, 2024). Rather, one requires the full joint (multivariate) distribution of both effect modifiers and purely prognostic covariates that are associated with the former. For non-collapsible measures such as the (log) odds ratio, the MTE depends on the full joint distribution of purely prognostic covariates, even if these are not associated with any effect modifiers or in the absence of treatment effect heterogeneity.
Table 3 presents a taxonomy of marginal estimands based on the heterogeneous illustrative models. Figure 2 indicates whether different estimands match for the heterogeneous illustrative models, across link functions and summary measures. The shaded squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent. In Table 3 and Figure 2, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
Table 3: Model-based marginal estimands for the heterogeneous illustrative models. For count outcomes and the log link, person-time is assumed constant, such that the log rate ratio is collapsible and can be interpreted as a log risk ratio.
Outcome Link function Summary measure Marginal estimand
Continuous Identity Mean difference Only depends on effect modifier means
Count Logarithmic Log risk ratio
Depends on the full joint distribution of effect modifiers and
purely prognostic covariates that are associated with the former
Binary Logit Log odds ratio
Depends on the full joint distribution of effect modifiers and
purely prognostic covariates
{logicpuzzle}
[…]
(a) Identity link, mean difference
{logicpuzzle}
[…]
(b) Logarithmic link, log rate ratio
{logicpuzzle}
[…]
(c) Logistic link, log odds ratio
Figure 2: Matrices indicating whether different estimands are equivalent for the heterogeneous illustrative models. The shaded squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
3.3 Non-linear (quadratic) heterogeneous CTEX
In the heterogeneous illustrative model in Section 3.2, the conditional outcome expectation on the linear predictor scale varies linearly with $X$, in both the active treatment and the control groups. Hence, the CTEX – the difference between the (transformed) conditional outcome expectations for each treatment – is also a linear function of $X$. In practice, such outcome expectations may vary non-linearly with the covariate in at least one of the treatment groups, in which case the CTEX is a non-linear function of $X$. We will now adopt another heterogeneous illustrative model, where the conditional outcome expectations on the linear predictor scale vary quadratically with $X$ in both treatment groups, such that the CTEX is a quadratic function of $X$.
The quadratic (heterogeneous) illustrative model specifies the conditional expectation of potential outcome $Y^{t}$ for treatment $T=t$ given $X$ as:
$E(Y^{t}\mid X)=g^{-1}\left(\beta_{0}+\beta_{1}X+\beta_{2}X^{2}+\beta_{T}t+\beta_{1T}Xt+\beta_{2T}X^{2}t\right).$ (7)
Again, $X$ is prognostic of outcome and a (conditional) treatment effect modifier on the linear predictor scale. In this case, the effect modification mechanism is more complex. Due to the second-order interaction – that is, the squared covariate-treatment product term – subjects with extreme values of $X$ experience lesser treatment efficacy, whereas subjects with moderate covariate values experience greater treatment efficacy, or vice versa (Mayer et al., 2017).
For a given link function, substituting Equation 7 into Equation 2 yields $\textnormal{CTEX}=\beta_{T}+\beta_{1T}x+\beta_{2T}x^{2}$ at $X=x$ on the linear predictor scale. That is, for the identity, logarithmic or logit link, the corresponding conditional mean difference, log risk ratio or log odds ratio, respectively, depends on the values of the covariate and the squared covariate. For a given link function, applying Equation 7 to Equation 3 yields $\textnormal{CTEM}=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}\bar{X}^{2}$, a quadratic function of the covariate mean.
[…]
What we seek to show with the quadratic illustrative model is that, unlike for the model in Section 3.2, the CTEM and the PACTE are not generally equivalent. Where the CTEX is constant or a linear function of $X$, it just so happens that the PACTE expression collapses to the CTEM on the linear predictor scale. Nevertheless, this is not the case where the CTEX is a non-linear function of $X$. Here, the interpretation of the CTEM is restricted to the covariate mean and no longer reflects a population-average conditional effect. Likewise, and contrary to suggestions in the literature (Remiro-Azócar et al., 2021a), the PACTE is a population-average estimand that does not necessarily represent the effect for subjects with the mean covariate values.
With respect to the marginal estimand, consider that $g(\cdot)$ is the identity link. On the mean difference scale, by the law of total expectation:
MTE $=E\left(Y^{1}\right)-E\left(Y^{0}\right)$
$=E_{X}\left[E\left(Y^{1}\mid X\right)\right]-E_{X}\left[E\left(Y^{0}\mid X\right)\right]$
$=E_{X}\left[\beta_{0}+\beta_{1}X+\beta_{2}X^{2}\right]+E_{X}\left[\beta_{T}+\beta_{1T}X+\beta_{2T}X^{2}\right]-E_{X}\left[\beta_{0}+\beta_{1}X+\beta_{2}X^{2}\right]$
$=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}E[X^{2}],$
such that the marginal treatment effect estimand coincides with the PACTE but not with the CTEM. Moreover, making use of the formula $Var(X)=E(X^{2})-\bar{X}^{2}$, yields the marginal mean difference $\textnormal{MTE}=\beta_{T}+\beta_{1T}\bar{X}+\beta_{2T}\left(\bar{X}^{2}+Var(X)\right)$. Consequently, and due to the direct collapsibility of the mean difference, the marginal estimand can be reduced to simple terms involving model coefficients and the mean and variance of the effect modifier $X$.
Conversely, as per Section 3.2, this is not the case for the marginal (log) risk ratio with $g(\cdot)=\log(\cdot)$, and for the marginal (log) odds ratio with $g(\cdot)=\operatorname{logit}(\cdot)$, for which the MTE on the corresponding linear predictor scale is not equivalent to the CTEM or to the PACTE. Again, for these outcome-generating models and summary measures, the MTE depends on the full distribution of covariate $X$ (not just its mean and its variance), and also the distribution of baseline risk, as prognostic factors may not even be known or measured. With multiple prognostic baseline covariates and in the presence of treatment effect heterogeneity, the MTE will generally depend on the full joint distribution of both effect modifiers and purely prognostic covariates (Remiro-Azócar, 2024).
Table 4 presents a taxonomy of marginal estimands based on the quadratic illustrative models. Figure 3 indicates whether different estimands match for the quadratic illustrative models, across link functions and summary measures. The shaded squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent. In Table 4 and Figure 3, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
Table 4: Model-based marginal estimands for the quadratic (heterogeneous) illustrative models. For count outcomes and the log link, person-time is assumed constant, such that the log rate ratio is collapsible and can be interpreted as a log risk ratio.
Outcome Link function Summary measure Marginal estimand
Continuous Identity Mean difference Depends on effect modifier means and variances
Count Logarithmic Log risk ratio
Depends on the full joint distribution of effect modifiers and
purely prognostic covariates that are associated with the former
Binary Logit Log odds ratio
Depends on the full joint distribution of effect modifiers and
purely prognostic covariates
{logicpuzzle}
[…]
(a) Identity link, mean difference
{logicpuzzle}
[…]
(b) Logarithmic link, log risk ratio
{logicpuzzle}
[…]
(c) Logistic link, log odds ratio
Figure 3: Matrices indicating whether different estimands are equivalent for the quadratic (heterogeneous) illustrative models. The shaded squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
4 The case of indirect treatment comparisons
We have clarified the interpretation of different marginal and conditional summary measures, considering different types of outcomes and outcome-generating mechanisms, with and without effect modification. In order to illustrate potential implications of estimand incompatibility for evidence synthesis, we consider the following two-study scenario, increasingly encountered in health technology assessment (Phillippo et al., 2016, 2018).
A manufacturer submitting evidence for reimbursement has conducted its own “index” RCT comparing the efficacy of novel active treatment $A$ versus control $C$. Another company has previously developed a competing active treatment and performed an RCT (the “competitor” trial) comparing the efficacy of its product, active treatment $B$, against control $C$. An indirect treatment comparison between $A$ and $B$, anchored relative to common comparator $C$, is required for reimbursement purposes. Typically, the manufacturer submitting the evidence has individual patient data (IPD) from the index RCT, but only aggregate-level summary data from the competitor RCT. For the latter trial, access to subject-level data is unavailable due to privacy and confidentiality concerns. This typically restricts the target population of the analysis to the competitor trial population.
Covariate adjustment methods – denoted “population adjustment” in this context – have been proposed to project the treatment effect for $A$ versus $C$ from the index trial to the competitor trial. A “fair” indirect comparison is then performed in the competitor trial population. Namely, the treatment effect for $A$ versus $B$ is estimated by combining treatment effect estimates for $A$ versus $C$ and for $B$ versus $C$, on the additive scale:
$\hat{\Delta}_{AB}^{(BC)}=\hat{\Delta}_{AC}^{(BC)}-\hat{\Delta}_{BC}^{(BC)},$ (8)
where $\hat{\Delta}_{tt^{\prime}}^{(BC)}$ is the effect estimate for $t$ versus $t^{\prime}$, with superscript $BC$ indexing the competitor trial population.
To produce the estimate $\hat{\Delta}_{AC}^{(BC)}$ in Equation 8, several population adjustment methods have been developed: (1) matching-adjusted indirect comparison (MAIC) (Signorovitch et al., 2010); (2) simulated treatment comparison (STC) (Ishak et al., 2015); (3) multilevel network meta-regression (ML-NMR) (Phillippo et al., 2020); (4) network meta-interpolation (NMI) (Harari et al., 2023); and (5) cross-network meta-regression (cross-NMR) (Hamza et al., 2023, 2024). An oft-overlooked aspect is that these estimate different summary measures, as outlined in Table 5. MAIC estimates the MTE; the “plug-in” or “mean-centering” STC version (Phillippo et al., 2016) estimates the CTEM, but simulation- or G-computation-based adaptations (Remiro-Azócar et al., 2022) estimate the MTE; ML-NMR can estimate the PACTE or the MTE (Phillippo et al., 2023); and NMI and cross-NMR estimate the CTEM.
Table 5: Summary measures estimated by different covariate adjustment approaches in the context of indirect comparisons.
Methodology Summary measure
Matching-adjusted indirect comparison MTE
Simulated treatment comparison (plug-in) CTEM
Simulated treatment comparison (G-computation) MTE
Multilevel network meta-regression PACTE and MTE
Network meta-interpolation CTEM
Cross-network meta-regression CTEM
In MAIC and the G-computation version of STC, $\hat{\Delta}_{AC}^{(BC)}$ is an estimate of the MTE. As such, the $\hat{\Delta}_{BC}^{(BC)}$ plugged into Equation 8 must estimate a compatible marginal summary measure. This is not problematic; a point estimate of the MTE is typically reported in RCT publications. Alternatively, “crude” unadjusted point estimates of the MTE can be derived from published aggregate-level outcome data, e.g., for binary outcomes, marginal risk differences, log risk ratios or log odds ratios can be calculated from event counts reported in two-by-two contingency tables.
The plug-in version of STC, NMI and cross-NMR plug in mean covariate values for the competitor trial population when estimating the covariate-adjusted treatment effect for $A$ versus $C$, such that $\hat{\Delta}_{AC}^{(BC)}$ is an estimate of the CTEM. However, any estimate $\hat{\Delta}_{BC}^{(BC)}$ that is available for $B$ versus $C$ is likely an incompatible marginal or conditional summary measure. RCTs often report estimates of conditional estimands, e.g., where the published effect estimate is the treatment coefficient of a multivariable regression of outcome on treatment and baseline covariates. For simplicity, our exposition in Section 2 and Section 3 has considered a single baseline covariate, and that conditional summary measures conditioning on said covariate. Nevertheless, conditioning on different sets of covariates leads to different conditional summary measures or estimands.
Trials with conditional estimates are likely to report effects conditioned on different covariate sets. These are not necessarily compatible across trials – e.g., for non-collapsible measures – even if the trials have identical target populations (Daniel et al., 2021). For a given target population, there are a multitude of PACTEs and CTEMs, one for every combination of baseline covariates considered for conditioning. For the plug-in STC and NMI, if conditional estimates are reported for the competitor trial, both $\hat{\Delta}_{AC}^{(BC)}$ and $\hat{\Delta}_{BC}^{(BC)}$ would have to condition on the same covariate set to guarantee estimand compatibility.
ML-NMR is not susceptible to estimand incompatibility issues. The multilevel regression allows evidence to be combined at the level of the CTEX, directly producing CTEX estimates for $A$ versus $B$. PACTE estimates for $A$ versus $B$ are obtained by averaging the CTEX on the linear predictor scale. The individual-level outcome model can be used to predict outcomes under treatments $A$ and $B$, on their natural scale. Averaging the predictions yields marginal outcome mean estimates under each treatment, and one can produce an MTE estimate for $A$ versus $B$ by contrasting the back-transformed marginal outcome means on the scale of interest (Phillippo et al., 2023).
Importantly, population-level decision-making requires population-average estimates that are relevant to the decision target population. The target population need not be represented by that of the competitor trial or may not be reflected by any of the studies in the evidence base. An advantage of ML-NMR is that it is not necessarily limited to produce treatment effect estimates in the competitor trial population, and can produce population-average marginal and conditional estimates in any target population of interest. This typically requires an additional identifying assumption to be made, which may be relaxed or assessed in larger evidence networks.(Phillippo et al., 2020, 2023).
5 Discussion and concluding remarks
While ICH E9 (R1) lacks explicit guidelines for estimands in evidence synthesis, it already warns of potential estimand incompatibilities, cautioning against naïve comparisons across trials without consideration of the estimand that is addressed in each.
We have clarified the interpretation of different marginal and conditional summary measures, and illustrated practical implications of estimand incompatibility for indirect treatment comparisons. Several novel covariate adjustment methodologies have been developed in this area, with different approaches targeting different summary measures. Nevertheless, the described estimand incompatibility issues may also apply to other areas of evidence synthesis, such as the naïve pooling of marginal and conditional estimates in pairwise – unadjusted or covariate-adjusted – meta-analyses of direct treatment comparisons (stat00513, 2016), or across indirect treatment comparisons or network meta-analyses that do not adjust for covariates (stat07909, 2014).
Estimand incompatibility issues are particularly concerning for evidence synthesis based on published summary effect measures, as different trials may report different summary measures. Full access to subject-level data for all trials is rarely available in practice. Nevertheless, this would be beneficial to resolve estimand incompatibility issues by allowing for an IPD (network) meta-regression, the “gold standard” evidence synthesis approach that ML-NMR seeks to generalize. Such approach can: (1) produce any desired marginal or conditional summary measure without incompatibility issues; (2) resolve heterogeneity due to population differences; and (3) perform treatment comparisons in any given target population of interest.
Many evidence syntheses of aggregate-level data are based on raw outcome data reported in each study arm, e.g., event counts and sample sizes, and thus combine marginal treatment effects across studies. However, purely prognostic covariates that are not conventionally described as effect modifiers – in the sense that they do not interact with treatment or induce treatment effect heterogeneity across subjects – can modify marginal treatment effects and compromise transportability, particularly for non-collapsible summary measures. This is generally under-appreciated in the evidence synthesis literature, with the exception of a recent article on network meta-analysis (Riley et al., 2023). Whether particular covariates can be classed as effect modifiers – thereby leading to departures from transportability, constancy or external validity – depends on the summary measure, and on whether it is marginal or conditional.
While the summary measure is at the heart of the estimand framework, as described by ICH E9 (R1), it is not an explicit component of frameworks such as PICO (Population, Intervention, Comparator, Outcome), typically used for scoping in evidence synthesis, and often a secondary consideration when postulating research questions in evidence synthesis. Ultimately, we seek to underscore the importance of the summary measure as a fundamental component of any clinical research question.
Within evidence synthesis, we can think of estimands at two different hierarchical levels: (1) at the trial level; and (2) at the meta-analytical level (Remiro-Azócar, 2022a). Both levels are important. There is a need to align estimands across different studies, as highlighted by this article. Nevertheless, there is also a need to align the estimand targeted at the meta-analytical level with a relevant clinical research question, motivating the shift towards more targeted meta-analyses that enhance applicability with respect to healthcare decision-making contexts (Ades et al., 2024; Remiro-Azócar & Gorst-Rasmussen, 2024).
[…]
Not available.
Related articles
stat00513 (2016), stat03728 (2014), stat05130 (2015), stat05152 (2015), stat07909 (2014).